It took 10 hours for me to get from Maryland to Denmark. How many seconds was that?

Answer by jim_thompson5910(35256)

1 minute = 60 seconds

60*1 minute = 60*60 seconds

60 minutes = 3600 seconds

1 hour = 3600 seconds

10*1 hour = 10*3600 seconds

10 hours = 36000 seconds